1. Economists have made estimates of the price elasticity of demand for a variety of goods and services.
  2. What is the price elasticity of demand and why is it important to estimate it? Provide your own examples to support your answer.

 

    ii.    The government commissioned a research firm, Super Consulting, to conduct a study on the market demand for cigarettes. The firm reported that the price elasticity of demand for cigarettes is about 0.4. If a pack of cigarettes costs $2 and the government wants to reduce smoking by 20 percent, by how much should the price increase using the midpoint method?


1
Expert's answer
2021-05-26T11:53:13-0400

i) Price elasticity of demand is the responsiveness of quantity demanded to change in price. That is, how quantity demanded changes with change in price.


If percentage change in quantity demanded is more than the percentage change in price then demand is said to be elastic. And value of elasticity is greater than 1. Luxury goods have elastic demand. Eg- expensive clothes, antiques etc.


ii) Price elasticity of demand is 0.4

Pack of cigarettes costs $2

and the government wants to reduce smoking by 20%

Price elasticity of demand is given by (% Change in Quantity demanded ÷ % Change in Price)

0.4 =( 20% ÷ % Change in price)

% Change in Price = (20% ÷0.4)

% Change in Price = 50%

Thus as demand for smoking declines price will have to increase by 50%

% Change in Price = "([\frac {New price } {Old Price}] - 1) × 100"


"50 = [ \frac {New Price} {\$ 2} - 1 ] × 100"


2 × 1.50

= $3


"[ ( \frac {\$3} {\$2} ) - 1 ] × 100"


= ( 1.50 - 1) × 100

= 0.50×100

= 50%

Thus the price will have to be increased by 50% to $3